OpenStudy (anonymous):

HELP! who can explain derivatives :( let f(x)= (x^3+3x+8)^2 F'(x)= F'(2)=

OpenStudy (anonymous):

use the chain rule get \[F'(x)=2(x^2+3x+8)(3x^2+3)\]

OpenStudy (anonymous):

I am very bad with derivatives. I missed the whole lesson de to being sick. can we start from the beginning?

OpenStudy (anonymous):

how do we use the chain rule?

OpenStudy (anonymous):

you have something squared the derivative of something squared is two times something, times the derivative of something

OpenStudy (anonymous):

for example, the derivative of \(\sin^2(x)\) is \(2\sin(x)\cos(x)\)

OpenStudy (anonymous):

the derivative of \(x+\sqrt{x}+\tan(x))^2\) is \[2(x+\sqrt{x}+\tan(x))(1+\frac{1}{2\sqrt{x}}+\sec^2(x))\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so how do we begin this problem?

OpenStudy (anonymous):

to find the derivative we need to place the exponent in front... correct?

OpenStudy (anonymous):

is the derivative for f'(x) = 2X^2+6X+16?